# DP-LET : An Efficient Spatio-Temporal Network Traffic Prediction Framework

### v1

@@ Abstract
Index Terms:
Traffic Prediction, Deep Learning, Transformer, Performance Evaluation.
@@ III-A Data Processing Module (Preprocessing Part)
In the preprocessing stage, the input spatio-temporal data ${X}\in\mathbb{R}^{M\times L}$is first denoised via TSVDR, where $M$ denotes the number of channels and $L$ denotes the input time horizon. Channel independence is applied to decouple the spatio-temporal data into $M$ univariate time series $\{X^{(i)}\in\mathbb{R}^{1\times L}\}_{i=1}^{M}$ to address challenges in spatial relationship modeling that arise after TSVDR denoising. This approach circumvents the pitfalls of imposing artificial spatial relationships while preserving critical temporal dynamics. Reversible instance normalization[15] and temporal patching mechanisms are incorporated to enhance data stationarity and capture multi-scale patterns, as illustrated in Algorithm 1.
@@ 14      $X^{(i)}_{\text{norm}}\leftarrow\frac{X^{(i)}-\mu_{i}}{\sigma_{i}}$;
[…]
@@ IV-B Experimental Result
$10^{5}$$10^{6}$$10^{7}$ineariTransformerPatchTSTTimeMixerDP-LETAutoformerInformerFedformerTimesNetCrossformerNumber of Parameters
Figure 2: Comparison of model parameter counts
Beyond predictive performance, we also evaluate the model complexity of each method by comparing the number of trainable parameters, as shown in Fig. 2. Among all Transformer-based approaches, DP-LET ranks third in terms of parameter count. Notably, it achieves a significant reduction in parameters compared to Transformer-based models that employ seasonal decomposition, such as Autoformer and FEDformer, nearly matching the model complexity of the MLP-based TimeMixer. This reduction highlights that our approach delivers high forecasting accuracy and maintains relatively low model complexity, effectively balancing efficiency and performance.

### v2

@@ III-A Data Processing Module (Preprocessing Part)
In the preprocessing stage, the input spatio-temporal data ${X}\in\mathbb{R}^{M\times L}$is first denoised via TSVDR, where $M$ denotes the number of channels and $L$ denotes the input time horizon. Channel independence is applied to decouple the spatio-temporal data into $M$ univariate time series $\{X^{(i)}\in\mathbb{R}^{1\times L}\}_{i=1}^{M}$ to address challenges in spatial relationship modeling that arise after TSVDR denoising. This approach circumvents the pitfalls of imposing artificial spatial relationships while preserving critical temporal dynamics. Reversible instance normalization and temporal patching mechanisms are incorporated to enhance data stationarity and capture multi-scale patterns, as illustrated in Algorithm 1.
@@ 14 $X^{(i)}_{\text{norm}}\leftarrow\frac{X^{(i)}-\mu_{i}}{\sigma_{i}}$;
15 $X^{(i)}_{\text{pad}}\leftarrow\bigl[X^{(i)}_{\text{norm}},\,0,\,0,\,\dots,\,0\bigr]$;
@@ IV-B Experimental Result
[…]
Figure 2: Efficiency analysis of DP-LET and baseline methods
Beyond predictive performance, we also evaluate the model complexity of each method by comparing the number of trainable parameters and time per training iteration, as shown in Fig. 2. Among Transformer-based approaches, DP-LET has the third-smallest parameter count. It substantially reduces both parameters and iteration time relative to seasonal-decomposition Transformers such as Autoformer and FEDformer, and its complexity is close to that of the MLP-based TimeMixer. These results demonstrate that our approach achieves high forecasting accuracy while maintaining a relatively light model footprint, effectively balancing efficiency and performance.
@@ V Conclusion
Acknowledgment
This work was partly supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number 23K28070.